OpenStudy (anonymous):

a car is traveling in a straight line and decelerating at a constant rate. after 3.2 seconds it has traveled 88m and slowed to a speed of 27 m/s what is the initial speed i got 52 m/s

OpenStudy (anonymous):

s=ut +1/2 *(a t^2) 88=u*3.2-(a/2)*3.2*3.2.......................1 v=u+at 27=u-a*3.2.....................2 multiply equation 2 with 1.6 and eliminate a and find out u (intial velocity) i got u=28

OpenStudy (anonymous):

i got 28 too